# Quench dynamics of the quantum XXZ chain with staggered interactions: Exact results and simulations on digital quantum computers

### v1

@@ I Introduction
Over the past decades, non-equilibrium dynamics of closed quantum systems has become an intensively studied topic, both theoretically and experimentally. Among experimental platforms, ultracold atomic systems offer an ideal setting for exploring such quantum dynamics [1]. More recently, rapid advances in programmable quantum devices based on various types of physical qubits have opened new avenues for probing the dynamics of quantum many-body systems.
Central questions in this field concern the nature of relaxation and the long-time behavior following a quantum quench, i.e. a sudden change in the Hamiltonian parameters. In homogeneous systems with extended eigenstates, local observables typically relax exponentially fast, and the stationary state depends on whether the underlying Hamiltonian is integrable or non-integrable. Non-integrable systems are expected to thermalize, reaching a stationary state described by a Gibbs ensemble with an effective temperature [2, 3, 5, 4]. On the other hand, integrable systems, characterized by an extensive number of conserved quantities, typically equilibrate to non-thermal stationary states described by a so-called generalized Gibbs ensemble [6, 7].
There also exist systems that fail to relax after a quench. Several mechanisms can lead to a lack of relaxation, including confinement [8, 9, 10], flat band [12], disorder [11], and many-body quantum scars [13]. Absence of relaxation is often reflected in the growth and spreading of entanglement; one scenario is persistent and long-lived entanglement oscillations even in the infinite-size limit.
@@
In addition, we perform numerical experiments on the IBM-Q quantum processors. In one experiment, we estimate the coefficients of the initial state in the Bell basis using the Hadamard test and reconstruct the time-dependent state from these coefficients to obtain dynamical entanglement entropies and the Loschmidt echo for various values $\Delta$. In a second experiment, we implement quantum circuits for the time evolution and measure the evolved states in randomly selected Pauli bases. Using classical post-processing of these measurement outcomes via statistical correlations and classical shadows [30, 31, 32], we estimate the second-order Rényi entanglement entropy and the Loschmidt echo.
@@ II Model and its non-equilibrium dynamics
For $\Delta=0$, the model reduces to the XX chain, which can be mapped to well-studied free-fermion systems [14, 15]. At $\Delta=1$, the system becomes the XXX chain: in the uniform case ($J=J^{\prime}$), it is exactly solvable via the Bethe ansatz [16], while in the dimerized case it exhibits the well-known spin-Peierls transition [17].
@@
where $\mathcal{N}=2^{2n}$. The expansion coefficients $a_{k}=\langle{\Phi_{k}}|{\Psi_{0}}\rangle$, given by scalar products between products of Bell states and singlet states, are all real. Thus, once the set of $\Delta$-independent coefficients $a_{k}$ is known, the time-dependent state at any $t$ under the postquench Hamiltonian with arbitrary anisotropy parameter $\Delta$ can be determined.
@@ II.2 Dynamical quantities
The Loschmidt echo, also known as the return probability, quantifies the overlap between the initial state $\left|{\Psi_{0}}\right\rangle$ and its time-evolved counterpart $\left|{\Psi_{0}(t)}\right\rangle$:
@@
The Loschmidt echo can be interpreted as the squared magnitude of an out-of-equilibrium analog of the partition function in equilibrium statistical mechanics. In the thermodynamic limit, zeros in the Loschmidt echo signal a dynamical phase transition, with singularities appearing in the return rate function [22, 23, 24]. Exact zeros of the Loschmidt echo can also occur in finite-size systems under certain conditions, as reported in recent studies [25, 26, 27]. In this work, we identify finite-size Loschmidt zeros in our quench experiments for specific anisotropy parameters.
@@ III.1.2 $n=3,5,\cdots$
To clarify this, Fig. 2 illustrates the initial singlet-product state (gray lines) and the couplings (red lines) in the postquench Hamiltonian for chains of length $n=2,4$ (Fig. 2(a) and (b)) and $n=3,5$ (Fig. 2 (c) and (d)), where the dashed line indicates the symmetric bipartition. In our setup, each initial entangled dimer located at $(i,i+1)$ can spread over three lattice spacings, covering sites from i1i-1 to $i+2$, during the time evolution [28, 21]. The EE of a half chain is the number of entangled states that connect sites inside the partition to sites outside. For any odd $n>2$, the initial state always contains an entangled dimer that spans the boundary between sites $i=n$ and $i=n+1$, contributing a time-independent unit to the EE; this is the constant term ’$1$’ in Eq. (33). In addition, at most two entangled states may cross the boundary between sites $i=1$ and $i=2n$, evolved from the singlets initially at $(1,2)$ and (2n1,2n)(2n-1,2n). This time-dependent part of the EE is analogous to the $n=2$ case. However, in the $n=2$ chain, the short chain length forces a single entangled state to cross both boundaries simultaneously, whereas for $n>2$ it can cross only one boundary. As a result, one unit of the EE in the $n=2$ case involves two couplings with a total strength $2J$ in the postquench Hamiltonian (red lines in Fig. 2), while for $n>2$ it involves only a single postquench coupling of strength $J$. This difference leads to a doubling of the entanglement oscillation period for odd $n>2$ compared with the $n=2$ case.
@@ III.2 Loschmidt echo
Figure 8: Time dependence of the Loschmidt echo [(a) and (c)] and the return rate function [(b) and (d)] with an irrational value of the anisotropy parameter $\Delta=2/(\sqrt{5}+1)$ for the XXZ chain of length $2n$ with even $n$ [upper panel (a) and (b)] and odd $n$ [upper panel (a) and (b)]. The coupling strength is fixed at $J=1$.
@@ IV.1 Hadamard test for $a_{k}$ estimation
Our first numerical experiment estimates the coefficients (amplitudes) $a_{k}=\langle{\Phi_{k}}|{\Psi_{0}}\rangle$ using the Hadamard test. The Hadamard test circuit, shown in Fig. 9, consists of two Hadamard gates $\bm{H}$ acting on a single ancilla qubit, which serves as the control qubit for a controlled-$\bm{U}$ gate. The unitary is defined as $\bm{U}=\bm{U}_{\Phi_{k}}^{\dagger}\bm{U}_{\Psi_{0}}$, where $\bm{U}_{\Psi_{0}}$ and $\bm{U}_{\Phi_{k}}$ generate the states $\left|{\Psi_{0}}\right\rangle$ and $\left|{\Phi_{k}}\right\rangle$, respectively, from a register of $N$ qubits initialized to $\left|{\bm{0}}\right\rangle=\left|{0}\right\rangle^{\otimes N}$:
@@
Figure 12: Simulation results (solid lines) for the half-chain entanglement entropies and the Loschmidt echo for a chain of length $N=4$ [Panel (a) -(c) ] and $N=6$ [Panel (d) and (e)] at various parameter $\Delta$, compared with exact solutions (dashed lines). The simulation data are based on the normalized coefficients $a_{k}$, estimated using the the Hadamard test on the quantum device ibm_fez for $N=4$ and on the noisy ibm simulator for $N=6$.
@@
Based on the selection rules described in Sec. II.1, we apply the Hadamard test exclusively to those active states $\left|{\Phi_{k^{*}}}\right\rangle$ that yield nonzero $a_{k^{*}}$. Results for $N=4$ and $N=6$, simulated on the IBM device ibm_fez and its noisy simulator, are shown in Fig. 11. For each data point, $N_{M}=2^{13}$ measurements per circuit for $N=4$ and $N_{M}=2^{14}$ for $N=6$ were performed. For $N=4$, only four active states contribute to the superposition: three with positive coefficients and one with a negative coefficient. The Hadamard test correctly reproduces the signs of these coefficients, though the measured magnitudes are systematically smaller than the exact values. For $N=6$, the number of active states increases to 16, and the accuracy of results from the real device declines significantly. This accuracy decline stems from two factors: the rapidly increasing circuit depth (exceeding $300$ for $N=6$) and the decreasing amplitude magnitudes ($\left|{a_{k^{*}}}\right|=0.25$ for $N=6$) for longer chains.
Figure 13: Simulation results (solid lines) for the half-chain Rényi entanglement entropy and the Loschmidt echo for a chain of $N=4$ with $\Delta=2/(\sqrt{5}+1)$. The simulation data are based on the normalized coefficients $a_{k}$, estimated using the the Hadamard test on ibm_fez.
@@ IV.2 Time evolution circuit
in terms of Pauli gates $\bm{X}$, $\bm{Y}$ and $\bm{Z}$. The quantum circuits implementing the three exponential operators with $ZZ$, XXXX, and $YY$ couplings are illustrated in Fig. 14(a), 14(b), and 14(c), respectively. In this setup, the exponential operators are related to each other through the basis transformation: $({\bm{u}^{X}})^{\dagger}\bm{Z}\bm{u}^{X}=\bm{X}$ and $({\bm{u}^{Y}})^{\dagger}\bm{Z}\bm{u}^{Y}=\bm{Y}$ with $\bm{u}^{X}=\bm{H}$ and $\bm{u}^{Y}=\bm{HS}^{\dagger}$.
@@ IV.3 Randomized measurements
The measurement protocol proceeds as follows: (i) Prepare the postquench state $\rho(t)$ of $N$ qubits at time $t$ using the time-evolution circuit. (ii) Apply a random unitary transformation, $\rho\rightarrow\bm{U}\rho\bm{U}^{\dagger}$, where $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ and each $\bm{u}_{i}$ is randomly drawn from $\{\bm{H},\bm{H}\bm{S}^{\dagger},\bm{I}\}$, thereby rotating the local measurement basis to a randomly chosen Pauli basis $\{X,Y,Z\}$. (iii) Measure each qubit in the standard $Z$ basis, obtaining a bitstring $\mathbf{b}=z_{1}\cdots z_{N}$ with $z_{i}=0,1$. For each fixed unitary $\bm{U}$, the measurement is repeated $N_{M}$ times. The full procedure is then repeated for $N_{U}$ independently sampled unitaries. The resulting $N_{N}\times N_{M}$ bitstrings and corresponding sets of single-qubit unitaries are stored for classical post-processing.
@@
for $m,\,m^{\prime}=1,\cdots,M$. We have used the following identity to efficiently compute the the trace of the product:
@@ V Summary and discussion
We have analytically and numerically investigated quantum quenches in the XXZ chain with alternating interactions in the flat-band limit. The quench protocol we considered interchanges the odd- and even-bond strengths of an initially fully dimerized chain. By formulating the problem in the Bell basis, we established selection rules identifying all Bell states with nonzero expansion coefficients in the time-dependent state for arbitrary even system sizes. For a chain of length $N$, all nonzero coefficients have equal magnitude $\left|{a_{k}}\right|=2^{-(N/2-1)}$. We further derived exact relations connecting the dynamical half-chain entanglement entropies of the $N=4$ system to those of any even system with $N>4$. From the exact results for the Loschmidt echo up to $N=16$, we identified a series of Loschmidt zeros occurring at critical times in certain finite chains. No such zeros were found in the isotropic XXX case ($\Delta=1$) or in chains whose length is not a multiple of four. Both the entanglement entropy and the Loschmidt echo exhibit persistent oscillations, reflecting the absence of relaxation in this flat-band system. For $N>4$, these observables are periodic with period with period $2q\pi$ when $J\Delta=p/q$ is an irreducible fraction.
@@
Several directions merit further exploration. Establishing a complete characterization of the system-size dependence of Loschmidt zeros and their corresponding critical times would provide deeper insight into finite-size effects in quench dynamics [25, 26, 27]. For quantum hardware implementations, shallow-circuit variants of the Hadamard test could enable amplitude estimation for intermediate system sizes [35, 36, 37]. Developing scalable, hardware-efficient protocols for measuring Loschmidt echoes and probing dynamical phase transitions remains an important challenge for using noisy intermediate-scale quantum devices in studies of non-equilibrium phenomena.
@@ Acknowledgements.
We acknowledge support from the IBM Q Hub at National Taiwan University, and the National Science and Technology Council (NSTC). Y.C.L. was supported by the NSTC under Grants No. NSTC 114-2119-M-007-013 and 113-2119-M-007-013; she is grateful to H.-C. Hsu and H.-C. Chang for useful discussions and previous collaboration. F.I. acknowledges travel support from the NSTC and National Chengchi University.

### v2

@@ I Introduction
Over the past decades, non-equilibrium dynamics of closed quantum systems has become an intensively studied topic, both theoretically and experimentally. Among experimental platforms, ultracold atomic systems offer an ideal setting for exploring such quantum dynamics [1, 2, 3, 4, 5]. More recently, rapid advances in programmable quantum devices based on various types of physical qubits have opened new avenues for probing the dynamics of quantum many-body systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
Central questions in this field concern the nature of relaxation and the long-time behavior following a quantum quench, i.e. a sudden change in the Hamiltonian parameters [17]. In homogeneous systems with extended eigenstates, local observables typically relax exponentially fast, and the stationary state depends on whether the underlying Hamiltonian is integrable or non-integrable. Non-integrable systems are expected to thermalize, reaching a stationary state described by a Gibbs ensemble with an effective temperature [18, 19, 20, 21]. On the other hand, integrable systems, characterized by an extensive number of conserved quantities, typically equilibrate to non-thermal stationary states described by a so-called generalized Gibbs ensemble [22, 23, 24, 25, 26, 27, 28, 29].
There also exist systems that fail to relax after a quench. Several mechanisms can lead to a lack of relaxation, including confinement [30, 31, 32], flat band [33, 34], disorder [35, 36], and many-body quantum scars [37, 38]. Absence of relaxation is often reflected in the growth and spreading of entanglement; one scenario is persistent and long-lived entanglement oscillations even in the infinite-size limit.
@@
In addition, we perform numerical experiments on the IBM-Q quantum processors. In one experiment, we estimate the coefficients of the initial state in the Bell basis using the Hadamard test [39] and reconstruct the time-dependent state from these coefficients to obtain dynamical entanglement entropies and the Loschmidt echo for various values $\Delta$. In a second experiment, we implement quantum circuits for the time evolution and measure the evolved states in randomly selected Pauli bases. Using classical post-processing of these measurement outcomes via statistical correlations and classical shadows [40, 41, 42, 43, 44], we estimate the second-order Rényi entanglement entropy and the Loschmidt echo.
@@ II Model and its non-equilibrium dynamics
For $\Delta=0$, the model reduces to the XX chain, which can be mapped to well-studied free-fermion systems [45, 46]. At $\Delta=1$, the system becomes the XXX chain: in the uniform case ($J=J^{\prime}$), it is exactly solvable via the Bethe ansatz [47], while in the dimerized case it exhibits the well-known spin-Peierls transition [48, 49].
@@
where $\mathcal{N}=2^{2n}$. The expansion coefficients $a_{k}=\langle{\Phi_{k}}|{\Psi_{0}}\rangle$, given by scalar products between products of Bell states and singlet states, are all real. Once the set of $\Delta$-independent coefficients $a_{k}$ is known, the time-dependent state at any $t$ under the postquench Hamiltonian with arbitrary anisotropy parameter $\Delta$ can be determined.
@@ II.2 Dynamical quantities
The Loschmidt echo, also known as the return probability, quantifies the overlap between the initial state $\left|{\Psi_{0}}\right\rangle$ and its time-evolved counterpart $\left|{\Psi_{0}(t)}\right\rangle$ [53, 54, 55, 56, 57, 58, 59]:
@@
The Loschmidt echo can be interpreted as the squared magnitude of an out-of-equilibrium analog of the partition function in equilibrium statistical mechanics. In the thermodynamic limit, zeros in the Loschmidt echo signal a dynamical phase transition, with singularities appearing in the return rate function [53, 54, 55, 56, 57, 58, 59]. Exact zeros of the Loschmidt echo can also occur in finite-size systems under certain conditions, as reported in recent studies [60, 61, 62]. In this work, we identify finite-size Loschmidt zeros in our quench experiments for specific anisotropy parameters.
@@ III.1.2 $n=3,5,\cdots$
To clarify this, Fig. 2 illustrates the initial singlet-product state (gray lines) and the couplings (red lines) in the postquench Hamiltonian for chains of length $n=2,4$ (Fig. 2(a) and (b)) and $n=3,5$ (Fig. 2 (c) and (d)), where the dashed line indicates the symmetric bipartition. In our setup, each initial entangled dimer located at $(i,i+1)$ can spread over three lattice spacings, covering sites from i1i-1 to $i+2$, during the time evolution [63, 13]. The EE of a half chain is the number of entangled states that connect sites inside the partition to sites outside. For any odd $n>2$, the initial state always contains an entangled dimer that crosses the boundary between sites $i=n$ and $i=n+1$, contributing a time-independent unit to the EE; this is the constant term ’$1$’ in Eq. (33). In addition, at most two entangled states may cross the boundary between sites $i=1$ and $i=2n$, evolved from the singlets initially at $(1,2)$ and (2n1,2n)(2n-1,2n). This time-dependent part of the EE is analogous to the $n=2$ case. However, in the $n=2$ chain, the short chain length forces a single entangled state to cross both boundaries simultaneously, whereas for $n>2$ it can cross only one boundary. As a result, one unit of the EE in the $n=2$ case involves two couplings with a total strength $2J$ in the postquench Hamiltonian (red lines in Fig. 2), while for $n>2$ it involves only a single postquench coupling of strength $J$. This difference leads to a doubling of the entanglement oscillation period for odd $n>2$ compared with the $n=2$ case.
@@ III.2 Loschmidt echo
Figure 8: Time dependence of the Loschmidt echo [(a) and (c)] and the return rate function [(b) and (d)] with an irrational value of the anisotropy parameter $\Delta=2/(\sqrt{5}+1)$ for the XXZ chain of length $2n$ with even $n$ [upper panel (a) and (b)] and odd $n$ [lower panel (c) and (d)]. The coupling strength is fixed at $J=1$.
@@ IV.1 Hadamard test for $a_{k}$ estimation
Our first numerical experiment estimates the coefficients (amplitudes) $a_{k}=\langle{\Phi_{k}}|{\Psi_{0}}\rangle$ using the Hadamard test [39]. The Hadamard test circuit, shown in Fig. 9, consists of two Hadamard gates $\bm{H}$ acting on a single ancilla qubit, which serves as the control qubit for a controlled-$\bm{U}$ gate. The unitary is defined as $\bm{U}=\bm{U}_{\Phi_{k}}^{\dagger}\bm{U}_{\Psi_{0}}$, where $\bm{U}_{\Psi_{0}}$ and $\bm{U}_{\Phi_{k}}$ generate the states $\left|{\Psi_{0}}\right\rangle$ and $\left|{\Phi_{k}}\right\rangle$, respectively, from a register of $N$ qubits initialized to $\left|{\bm{0}}\right\rangle=\left|{0}\right\rangle^{\otimes N}$:
@@
Figure 12: Simulation results (solid lines) for the half-chain entanglement entropies and the Loschmidt echo for a chain of length $N=4$ [Panel (a) -(c) ] and $N=6$ [Panel (d) and (e)] at various parameter $\Delta$, compared with exact solutions (dashed lines). The simulation data are based on the normalized coefficients $a_{k}$, estimated using the Hadamard test on the quantum device ibm_fez for $N=4$ and on the noisy ibm simulator for $N=6$.
@@
Based on the selection rules described in Sec. II.1, we apply the Hadamard test exclusively to those active states $\left|{\Phi_{k^{*}}}\right\rangle$ that yield nonzero $a_{k^{*}}$. Results for $N=4$ and $N=6$, simulated on the IBM device ibm_fez and its noisy simulator, are shown in Fig. 11. For each data point, $N_{M}=2^{13}$ measurements for $N=4$ and $N_{M}=2^{14}$ for $N=6$ were performed. For $N=4$, only four active states contribute to the superposition: three with positive coefficients and one with a negative coefficient. The Hadamard test correctly reproduces the signs of these coefficients, though the measured magnitudes are systematically smaller than the exact values. For $N=6$, the number of active states increases to 16, and the accuracy of results from the real device declines significantly. This accuracy decline stems from two factors: the rapidly increasing circuit depth (exceeding $300$ for $N=6$) and the decreasing amplitude magnitudes ($\left|{a_{k^{*}}}\right|=0.25$ for $N=6$) for longer chains.
Figure 13: Simulation results (solid lines) for the half-chain Rényi entanglement entropy and the Loschmidt echo for a chain of $N=4$ with $\Delta=2/(\sqrt{5}+1)$ and $\Delta=(\sqrt{5}+1)/2$. The simulation data are based on the normalized coefficients $a_{k}$, estimated using the Hadamard test on ibm_fez.
@@ IV.2 Time evolution circuit
in terms of Pauli gates $\bm{X}$, $\bm{Y}$ and $\bm{Z}$. The quantum circuits implementing the three exponential operators with $ZZ$, XXXX, and $YY$ couplings are illustrated in Fig. 14(a), 14(b), and 14(c), respectively. In this setup, the exponential operators are related to each other through the basis transformation: $({\bm{u}^{X}})^{\dagger}\bm{Z}\bm{u}^{X}=\bm{X}$ and $({\bm{u}^{Y}})^{\dagger}\bm{Z}\bm{u}^{Y}=\bm{Y}$ with $\bm{u}^{X}=\bm{H}$ and $\bm{u}^{Y}=\bm{HS}^{\dagger}$, where $\bm{S}$ is a phase shift gate which rotates a state vector a $\pi/2$ radian around the $Z$-axis.
@@ IV.3 Randomized measurements
The measurement protocol proceeds as follows: (i) Prepare the postquench state $\rho(t)$ of $N$ qubits at time $t$ using the time-evolution circuit. (ii) Apply a random unitary transformation, $\rho\rightarrow\bm{U}\rho\bm{U}^{\dagger}$, where $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ and each $\bm{u}_{i}$ is randomly drawn from $\{\bm{H},\bm{H}\bm{S}^{\dagger},\bm{I}\}$, thereby rotating the local measurement basis to a randomly chosen Pauli basis $\{X,Y,Z\}$. (iii) Measure each qubit in the standard $Z$ basis, obtaining a bitstring $\mathbf{b}=z_{1}\cdots z_{N}$ with $z_{i}=0,1$. For each fixed unitary $\bm{U}$, the measurement is repeated $N_{M}$ times. The full procedure is then repeated for $N_{U}$ independently sampled unitaries. The resulting $N_{U}\times N_{M}$ bitstrings and corresponding sets of single-qubit unitaries are stored for classical post-processing.
@@
for $m,\,m^{\prime}=1,\cdots,M$. We have used the following identity to efficiently compute the trace of the product:
@@ V Summary and discussion
We have analytically and numerically investigated quantum quenches in the XXZ chain with alternating interactions in the flat-band limit. The quench protocol we considered interchanges the odd- and even-bond strengths of an initially fully dimerized chain. By formulating the problem in the Bell basis, we established selection rules identifying all Bell states with nonzero expansion coefficients in the time-dependent state for arbitrary even system sizes. For a chain of length $N$, all nonzero coefficients have equal magnitude $\left|{a_{k}}\right|=2^{-(N/2-1)}$. We further derived exact relations connecting the dynamical half-chain entanglement entropies of the $N=4$ system to those of any even system with $N>4$. From the exact results for the Loschmidt echo up to $N=16$, we identified a series of Loschmidt zeros occurring at critical times in certain finite chains. No such zeros were found in the isotropic XXX case ($\Delta=1$) or in chains whose length is not a multiple of four. Both the entanglement entropy and the Loschmidt echo exhibit persistent oscillations, reflecting the absence of relaxation in this flat-band system. For $N>4$, these observables are periodic with period $2q\pi$ when $J\Delta=p/q$ is an irreducible fraction.
@@
Several directions merit further exploration. Establishing a complete characterization of the system-size dependence of Loschmidt zeros and their corresponding critical times would provide deeper insight into finite-size effects in quench dynamics [60, 61, 62]. Finite-rate quenches, in which systems’ parameters change gradually over a finite time, also present interesting avenues for extending this work [65, 66, 67, 68]. For quantum hardware implementations, shallow-circuit variants of the Hadamard test could enable amplitude estimation for intermediate system sizes [69, 70, 71, 72, 73, 74]. Developing scalable, hardware-efficient protocols for measuring Loschmidt echoes and probing dynamical phase transitions remains an important challenge for using noisy intermediate-scale quantum devices in studies of non-equilibrium phenomena.
@@ Acknowledgements.
We acknowledge support from the IBM Q Hub at National Taiwan University, and the National Science and Technology Council (NSTC). Y.C.L. was supported by the NSTC under Grants No. NSTC 114-2119-M-007-013 and 113-2119-M-007-013; she is grateful to H.-C. Hsu and H.-C. Chang for useful discussions and previous collaboration. F.I. was supported by the Hungarian Scientific Research Fund under Grant No. K146736 and by the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary; he also acknowledges travel support from the NSTC and National Chengchi University.